in a survey of 100 students, 56 wrote the Maths exams, 23 wrote psychology and 21 wrote the science exam. 12 wrote both maths and psychology exams, 9 write the maths and science exams and 6 wrote both psychology and science exams. 5 students wrote neither. determine how many students wrote all three exams.

Search Results
Web results
1
Expert's answer
2020-03-18T18:09:20-0400

We use the theory of inclusion and exclusion here.

Let "A,B,C" denote the events of writing the Mathematics, Psychology and Science Exams respectively.


Let "U" denote the universal set, so "N(U)=100" .


Now, we have the following information:

"N(A)=56\\

N(B)=23\\

N(C)=21\\

N(A\cap B)=12\\

N(A\cap C)=9\\

N(B\cap C)=6\\

N(\bar{A}\cap\bar{B}\cap\bar{C})=5"

From the last one, using the fact that "\bar{A}\cap\bar{B}\cap\bar{C}=(A\cup B\cup C)^C" , we have "N(A\cup B\cup C)=95"

We need "N(A\cap B\cap C)"

Using inclusion-exclusion principle, we have

"N(A\cup B\cup C)=N(A)+N(B)+N(C)\\\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-N(A\cap B)-N(A\cap C)-N(B\cap C)\\\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+N(A\cap B\cap C)\\
\implies N(A\cap B\cap C)=N(A\cup B\cup C)-(N(A)+N(B)+N(C))-(-N(A\cap B)-N(A\cap C)-N(B\cap C))\\\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\\
=95-56-23-21+12+9+6=22"